If the MPS is 0.4 and gov taxes decrease by 50 thousand, what is the change in Y?
Answer #1

Answer : Multiplier =1/MPS =1/0.4 = 2.5

Change in GDP = Multiplier * Change in government spending =2.5*50,000 =1,25,000